OpenStudy (anonymous):

Prove that \[\sqrt{5}\] is irrational.........

OpenStudy (anonymous):

I believe we would use a proof by contradiction.

OpenStudy (anonymous):

similar to the one with \[\sqrt2 \]right? :)

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

Could you do that for me?

OpenStudy (anonymous):

we will do that together...:)

OpenStudy (anonymous):

Of course... hehe

OpenStudy (anonymous):

suppose that \(\sqrt{5}\) is a rational number so we can write \[\sqrt{5}=\frac{p}{q}\ \ \ \ \ \gcd(p,q)=1 \]square both sides\[p^2=5q^2\]

OpenStudy (anonymous):

Yes.. I know it till now

OpenStudy (anonymous):

now suppose \(q\) is an even number then \(p\) will be even and this is impossible.

OpenStudy (anonymous):

then let \(q\) to be an odd number...\(p^2\) divides \(5\) so does \(p\) ... so \(p=5k\) plug in the equation \(25k^2=5q^2\) and \(q^2=5k^2\) so \(q^2\) divides \(5\) so does \(q\) and this is a contradiction

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw :)